OpenStudy (anonymous):

a girl has 240g of water at 50 degree centigrade .how many grams of ice at zero degree centigrade are needed to cool the water to zero degree centigrade a)100g b)140g c)150g d)58g

OpenStudy (anonymous):

@ganeshie8 @ParthKohli @e.mccormick @Luigi0210 @mathslover @matricked @paki

OpenStudy (anonymous):

use heat lost by water = mass of ice * Latent heat of ice

OpenStudy (anonymous):

in CGS 240 * 1*(50-0) = m* 80 m =150g

OpenStudy (anonymous):

Q =m cp delta T this formula ya used for second one ???

OpenStudy (anonymous):

H = mass * specific Latent heat H= M*L (symbols have standard meaning )

OpenStudy (anonymous):

ok got it . Thank you so much ^_^

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

@matricked Can ya help me in physics ?